OpenStudy (clairelizzz):

Solve the triangle A = 33°, a = 20, b = 13

OpenStudy (anonymous):

what type of triangle our we looking at?

OpenStudy (nincompoop):



OpenStudy (nincompoop):

are you familiar with laws of sine?

OpenStudy (nincompoop):



OpenStudy (nincompoop):

\(a/sinA = b/sinB = c/sinC \) or \(sinA/a = sinB/b = sinC/c\)

OpenStudy (nincompoop):

RESPOND

OpenStudy (nincompoop):

\(\large\frac{sinB}{13} = \frac{sin(33)}{20}\) make sure your calculator is set properly